比較文化演習 2 単位 (選択) 3 年 (前期), 4 年 (前期)

比較文化演習座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部

【授業目的】西洋・日本・イスラームにおける価値観・人権概念の比較分析，現代国際社会における多様な民主化理論や文化衝突，平和の構築という課題の重要性と戦争と平和の関係について，基本的に英語を用いて研究活動を行っていきます．

【授業概要】イスラーム思想を基本として，国際交流や国際問題，メディア報道など，私たちの身近にある異文化コミュニケーション現象を実例として取り上げ，ディスカッション・プレゼンテーション能力を高めていきます．英語を一つの言語として徹底理解することによって，さらに新たな異文化・言語の理解を促し，複眼的に文化を捉え，現代社会における多言語・多文化共生について考えていきましょう．

【キーワード】イスラーム，国際政治【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】イスラーム思想におけるイデオロギーや文化等を修める【授業計画】

1. Seminar General Guidance

2. An intriduction to Islamic Thoughts & Arabic Language1

3. An intriduction to Islamic Thoughts & Arabic Language 2

4. An intriduction to Islamic Thoughts & Arabic Language 3

5. An intriduction to Islamic Thoughts & Arabic Language 4

6. An intriduction to Islamic Thoughts & Arabic Language 5

7. An intriduction to Islamic Thoughts & Arabic Language 6

8. An intriduction to Islamic Thoughts & Arabic Language 7

9. An intriduction to Islamic Thoughts & Arabic Language 8

10. An intriduction to Islamic Thoughts & Arabic Language 9

11. An intriduction to Islamic Thoughts & Arabic Language10

12. An intriduction to Islamic Thoughts & Arabic Language11

13. An intriduction to Islamic Thoughts & Arabic Language12

14. An intriduction to Islamic Thoughts & Arabic Language13

15. An intriduction to Islamic Thoughts & Arabic Language14

16. Review 【成績評価】毎回の積極的な参加による【再試験】無し【教科書】座喜純，『イスラーム原理主義の「道しるべ」―発禁 “アルカイダの教

本” 全訳+解説』【参考書】座喜純，『常態のイスラーム―いまだ照らされていない世界』【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220149 【連絡先】

⇒ 座喜 . (オフィスアワー: Contact: jzaki@ias.tokushima-u.ac.jp)

比較文化演習 2 units (selection) 3rd-year(1st semester), 4th-year(1st semester)

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences

Target⟩ This seminar formed to give deep analyze of Islam as an ideology from

the approach of International Politics. Outline⟩ There will be practices in Islamic Thoughts, comparative studies on Arab

and Middle East Society and Culture. �������⟩ Islam, International Relation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Building a strong background reguarding Islam Thoughts and Islamic

Ideology

Schedule⟩

1. Seminar General Guidance

2. An intriduction to Islamic Thoughts & Arabic Language1

3. An intriduction to Islamic Thoughts & Arabic Language 2

4. An intriduction to Islamic Thoughts & Arabic Language 3

5. An intriduction to Islamic Thoughts & Arabic Language 4

6. An intriduction to Islamic Thoughts & Arabic Language 5

7. An intriduction to Islamic Thoughts & Arabic Language 6

8. An intriduction to Islamic Thoughts & Arabic Language 7

9. An intriduction to Islamic Thoughts & Arabic Language 8

10. An intriduction to Islamic Thoughts & Arabic Language 9

11. An intriduction to Islamic Thoughts & Arabic Language10

12. An intriduction to Islamic Thoughts & Arabic Language11

13. An intriduction to Islamic Thoughts & Arabic Language12

14. An intriduction to Islamic Thoughts & Arabic Language13

15. An intriduction to Islamic Thoughts & Arabic Language14

16. Review Evaluation Criteria⟩ weekly attendance, active participation and homework Re-evaluation⟩ no makeup exam to this class Textbook⟩ 座喜純，『イスラーム原理主義の「道しるべ」―発禁 “アルカイダの

教本” 全訳+解説』 Reference⟩ 座喜純，『常態のイスラーム―いまだ照らされていない世界』 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220149 Contact⟩ ⇒ Zaki . (Office Hour: Contact: jzaki@ias.tokushima-u.ac.jp)